Text Size

ISS Daily Summary Report - 05/20/13
May 20, 2013
 
International Space Station (ISS) Management Center (IMC) Daily Summary Report
Increment 36 Operations – Stage 36-3
05/20/13 4:00 pm Central Daylight Time (CDT)
140/21:00 Greenwich Mean Time (GMT)


New Status:
Multi-user Droplet Combustion Apparatus (MDCA) Hardware Replacement: On May 16, Combustion Integration Rack (CIR) experienced a failed igniter in the combustion chamber. Further test points could not be accomplished until it was replaced. Today, FE-3 Cassidy reviewed procedures, gathered hardware, accessed the CIR Combustion Chamber and removed the MDCA Chamber Insert Assembly (CIA) from the Chamber. He then replaced the MDCA Fuel Reservoir in Location 1 and both MDCA igniter tips and re-installed the MDCA CIA in the Chamber.

Treadmill with Vibration Isolation System (TVIS) Removal: CDR Vinogradov and FE-2 Misurkin removed TVIS and temporarily stowed it in the KAYUTA (Russian sleep quarters) in preparation for the new Russian Treadmill (BD-2) installation and checkout over the next 3 days. Crew is waiting until BD-2 is installed and operational before disassembling TVIS.

Removed Photo Voltaic Thermal Control System (PVTCS) PFCS Status: The Pump Flow Control System (PFCS) that was removed during US Extravehicular Activity (EVA) 21 on May 11, 2013, continued to show a steady leak rate after the EVA. Over the weekend the removed PFCS accumulator quantity fell to 0%. Ground controllers will uplink updated Pre-Positioned Load (PPLs) later today to allow PVTCS automated safing to be re-enabled on Tuesday. Ground controllers expect to reactivate Channel 2B Electrical Power System (EPS) hardware on Wednesday.

ISS Conjunction Threat: The ground was preparing for a possible ISS Debris Avoidance Maneuver (DAM) in response to a conjunction threat from object Kondor-E 1 Observation Satellite. Last night, the Probability of Collision (PC) was in the red threshold. At this time, the trend has gone to green and ground teams are standing down on previous DAM planning.

Today’s Planned Tasks
· RGN-REC TK R&R [Complete]
· Deactivating TVIS power circuit breaker [Complete]
· TVIS IFM Procedure Review [Complete]
· TVIS Tool Gathering [Complete]
· Cover preparation to protect the TVIS pit r/g 2396 [Complete]
· CIR worksite preparation [Complete]
· TVIS - Removal from SM Pit. [Complete]
· CIR Hardware Gathering [Complete]
· TVIS - Disassembly of Stabilizers, SLD/SPD Assemblies. [Complete]
· CIR Doors Open [Complete]
· CIR Front Cover Open [Complete]
· MDCA – Hardware Replacement [Complete]
· Temp Stow TVIS Elements in ТКГ 418 (DC1). Temp Stow in ТКГ 418 across plane I and IV [Complete]
· CIR - Seal Replacement [Complete]
· CIR Front Cover Close [Complete]
· CIR Doors Close [Complete]
· Test 1 of PCE, PCE antenna feeder (АФУ), ATV control panel commands. [Complete]
· CIR - Hardware Teardown [Complete]
· Laptop SSC 1 and 6 - Reconnect, Part 1 [Complete]
· TVIS - Frame Removal. [Complete]
· Temp Stow TVIS Elements in ТКГ 418 (DC1). Temp Stow in ТКГ 418 across plane I and IV [Complete]
· Installation of protective cover on TVIS pit / r/g 2396 [Complete]
· CIR Hardware Stow [Complete]
· Laptop SSC 7 - Wired Reconfig

Task List
· РСПИ (Data Transmission Radio Link). Photo/Video Data Prep on RSS2 for Downlink via РСПИ / r/g 2003
· Roskosmos report
· URAGAN
· ECON-M.
· MSG Logistics Drawer Audit [Complete]
· SSC Personal Data Copy
· WHC KTO replace
· WHC EDV R&R [Complete]
· CEVIS Pedal Clocking
· ARED Slide Tether
· Laptop Cleanup Part 5
· N2 Spare Tools Audit
· CWCI sort [Complete]
· D3S Sensor Clean
· Trash D2XS Camera
· P/TV Audit
· IMV Flow Measure

Ground
· Nominal Commanding

Three-Day Look Ahead:
Tuesday, May 21: BD-2 Install, Emergency OBT, CFE Hardware Setup, Fundoscope Exam
Wednesday, May 22: BD-2 Install, CFE ICF9 Test, Lab video encoder test
Thursday, May 23: BD-2 Install, SPHERES
Meeting Schedule:
Meeting
Day
Date
Houston Time (CDT)
IMMT
Thursday
May 23
8:00 a.m.